# Identity Inference from CLIP Models using Only Textual Data

### v1

@@ Abstract
The widespread usage of large-scale multimodal models like CLIP has heightened concerns about the leakage of personally identifiable information (PII). Existing methods for identity inference in CLIP models, i.e., to detect the presence of a person’s PII used for training a CLIP model, require querying the model with full PII, including textual descriptions of the person and corresponding images (e.g., the name and the face photo of the person). However, this may lead to potential privacy breach of the image, as it may have not been seen by the target model yet. Additionally, traditional membership inference attacks (MIAs) train shadow models to mimic the behaviors of the target model, which incurs high computational costs, especially for large CLIP models. To address these challenges, we propose a textual unimodal detector (TUNI) in CLIP models, a novel method for ID inference that 1) queries the target model with only text data; and 2) does not require training shadow models. Firstly, we develop a feature extraction algorithm, guided by the CLIP model, to extract features from a text description. TUNI starts with randomly generating textual gibberish that were clearly not utilized for training, and leverages their feature vectors to train a system of anomaly detectors. During inference, the feature vector of each test text is fed into the anomaly detectors to determine if the person’s PII is in the training set (abnormal) or not (normal). Moreover, TUNI can be further strengthened integrating real images associated with the tested individuals, if available at the detector. Extensive experiments of TUNI across various CLIP model architectures and datasets demonstrate its superior performance over baselines, albeit with only text data.
@@ 1 Introduction
Recent years have witnessed a rapid development of large-scale multimodal models, such as Contrastive Language–Image Pre-training (CLIP) (Radford et al., 2021). These models synthesize information across different modalities, particularly text and images, facilitating applications from automated image generation to sophisticated visual question answering systems. Despite their potential, these models pose significant privacy risks (Inan et al., 2021; Carlini et al., 2021; Leino and Fredrikson, 2020; Fredrikson et al., 2015; Rigaki and Garcia, 2023; Helbling et al., 2023; Rahman et al., 2024; Rahman, 2023) as the vast datasets used for training often contain personally identifiable information (PII) (Schwartz and Solove, 2011; Abadi et al., 2016; Bonawitz et al., 2017), raising concerns (Xi et al., 2024) about PII leakage and misuse (Hu et al., 2023; Yin et al., 2021). Therefore, it is extremely important to develop tools to detect potential PII leakage from CLIP models. Specially, as the first step, we would like to address the identity inference problem, i.e., to determine if the PII of a particular person was used in training of a target CLIP model.
Figure 1: Identity inference in a CLIP model with only textual data. In this example, a third-party detector, delegated by the owner of PII, queries the target CLIP model with only textual description of the person.
@@
To address these limitations, we propose a textual unimodel detector (TUNI) for identity inference in CLIP models, which queries the target model with only text information during inference. Specifically, we first propose a feature extractor, which maps a textual description to a feature vector though image optimization guided by the CLIP model; then, we randomly generate a large amount of textual gibberish, which we know do not match any textual descriptions in the training dataset. As shown in Figure 2, we make the key observation that the feature distributions of textual gibberish and member samples in the training set are well distinguishable.
@@ 2.1 Privacy Leakage in CLIP Models
The CLIP model exemplifies modern multimodal innovation by integrating an image encoder and a text encoder into its architecture (Radford et al., 2021). These encoders transform inputs into a shared embedding space, enabling effective measurement of semantic similarity (Ramesh et al., 2022). Despite the significant advances and expansive applicability of CLIP models, the vast and diverse datasets utilized for training such models could potentially include sensitive information, raising concerns about privacy leakage (Hu et al., 2022b). Various inference attacks, including model stealing  (Dziedzic et al., 2022; Liu et al., 2022; Wu et al., 2022), knowledge stealing (Liang et al., 2022), data stealing (He and Zhang, 2021), and membership inference attacks (Liu et al., 2021; Ko et al., 2023), have been developed for CLIP models, exposing potential vulnerability in privacy leakage. These privacy concerns underscore the necessity for developing robust defense mechanisms to safeguard sensitive information in CLIP models (Golatkar et al., 2022; Jia et al., 2023; Huang et al., 2023).
@@ 2.2 Personally Identifiable Information and Leakage Issues
Personally Identifiable Information (PII) is defined as any data that can either independently or when combined with other information, identify an individual. Training Large Language Models (LLMs) often utilizes publicly accessible datasets, which may inadvertently contain PII. This elevates the risk of data breaches that could compromise individual privacy and entail severe legal and reputational consequences for the deploying entities (Lukas et al., 2023; Abadi et al., 2016; Bonawitz et al., 2017; Rahman et al., 2020; Shamshad et al., 2023). Various attacks have been developed to reveal PII from LLMs. A method is proposed in (Panda et al., 2024) to steal private information from LLMs, via crafting specific queries to GPT-4 that can reveal sensitive data by appending a secret suffix to the generated text; Zhang et al. introduced the ETHICIST method for targeted training data extraction, through loss smoothed soft prompting and calibrated confidence estimation, significantly improving extraction performance on public benchmarks (Zhang et al., 2023); Carlini et al. also studied training data extraction from LLMs, emphasizing the predictive capability of attacks given a prefix (Carlini et al., 2021); ProPILE, proposed in (Kim et al., 2024), probes privacy leakage in LLMs, by assessing the leakage risk of PII included in the publicly available Pile dataset; Inan et al. investigated the risks associated with membership inference attacks using a Reddit dataset, further emphasizing the persistent threat of PII leakage in various data environments (Inan et al., 2021).
@@ 3.1 Problem Setup and Threat Model
Detector’s Knowledge and Capability. The detector operates with black-box access to $M$, i.e., it can query $M$ and observe the outputs, but does not know the model architecture of $M$, the parameter values, or the training algorithms. For the target textual description $t$, depending on the application scenarios, the detector may or may not have actual images corresponding to $t$. Nevertheless, in the case where the detector knows corresponding images, due to privacy concerns, it cannot include them in the queries to $M$. The detector cannot modify $M$ or access its internal state.
@@ 3.2 TUNI: Textual Unimodal Detector for ID Inference
Algorithm 1 CLIP-guided Feature Extraction.
@@
8:         $v_{x_{i}}\leftarrow M(x_{i})$triangleright Obtain image embedding from $M$
9:         $x_{i+1}\leftarrow\arg\max_{x_{i}}\frac{v_{t}\cdot v_{x_{i}}}{||v_{t}||\,||v_{x_{i}}||}$triangleright Update image to maximize cosine similarity
@@
11:     $S_{i}\leftarrow\frac{v_{t}\cdot v_{x_{m}}}{||v_{t}||\,||v_{x_{m}}||}$triangleright Optimized similarity for epoch $i$
12:     $\mathcal{S}\leftarrow\mathcal{S}\cup{S_{i}}$, $\mathcal{V}\leftarrow\mathcal{V}\cup v_{x_{m}}$
@@
16:$D\leftarrow\sqrt{\frac{1}{n}\sum_{V\in\mathcal{V}}||v-\bar{v}||^{2}}$
@@
Specifically, for each test text $t$, the detector is equipped with a set of $c$ real images $\{x_{\mathrm{real}}^{1},x_{\mathrm{real}}^{2},\ldots,x_{\mathrm{real}}^{c}\}$. Similar to the feature extraction process in Algorithm 1, over $k$ epochs with independent initializations, $k$ optimized images $\{x_{\mathrm{opt}}^{1},x_{\mathrm{opt}}^{2},\ldots,x_{\mathrm{opt}}^{k}\}$ for $t$ are obtained under the guidance of the CLIP model. Then, we apply a pretrained feature extraction model $F$ (e.g., DeepFace (Taigman et al., 2014) for face images) to the real and optimized images to obtain real embeddings $\{v_{\mathrm{real}}^{1},v_{\mathrm{real}}^{2},\ldots,v_{\mathrm{real}}^{c}\}$ and optimized embeddings $\{v_{\mathrm{opt}}^{1},v_{\mathrm{opt}}^{2},\ldots,v_{\mathrm{opt}}^{k}\}$. Finally, we compute the average pair-wise $\ell_{2}$ distance between the real and optimized embeddings, denoted by $R$, over $c\cdot k$ pairs, and use $R$ as an additional feature of the text $t$.
4 Evaluations
4.1 Setup
Dataset Construction. The datasets for training and ID inference are constructed from three datasets: LAION-5B (Schuhmann et al., 2022), Conceptual Captions 3M (CC3M) (Changpinyo et al., 2021), and FaceScrub (Kemelmacher-Shlizerman et al., 2016). Specifically, 200 celebrities—100 for training and 100 for validation, with their face photos accompanied by labels containing their names are selected from the FaceScrub dataset; then these data samples are augmented by additional photos of the selected celebrities found in LAION-5B, such that each person has multiple photos; finally these augmented data points are mixed with the CC3M dataset to form the training set of the CLIP model. By doing this, we have the ground truth on which people are in the training set and which are not. In our experiments, we construct two datasets, one with a single photo for each person, and another with 75 photos for each person. A more detailed description of the datasets is given in Appendix A.
Baselines. Current ID inference detection methods for CLIP models typically require the detector to query the target model with corresponding real images. Most MIAs involve training shadow models and related methods like shadow encoders  (Liu et al., 2021), which can be particularly costly for large-scale multimodal models. We empirically compare the performance of TUNI with the following SOTA inference methods, which both avoid using shadow models, but still require submitting both text and image to the target CLIP model for inference.
  • Identity Inference Attack (IDIA) (Hintersdorf et al., 2022) detects with a list of 1000 names to choose from and 30 real photos for a tested person. In IDIA, the attacker (detector) selects candidate names as prompt templates, and predicts names for each image and prompt. Then the attacker generates multiple queries by replacing the name and analyzes the model’s responses to calculate attack scores. Once the correct name is predicted, it’s inferred that the target individual is in the training data. We compare IDIA using 3 photos for each test sample with TUNI using only text.
4.2 Results
Table 1: Samples of randomly generated gibberish.
+̱7IKXb2Y FR!pnI<5xS euiT_;yw/ WN@D3Ô;(! Fi\.b"<.*B meeUx}jK"1
je|%5(s=G\_ ?Ŵ<E{Dvmz hqf- =j<q5 AAY5v1RWjR 2y?r"-r84{ dP[H:FR:B.
#lEZ0yrZ5ig ’2_:6[jiOa X*|<tFx|4/ {+W"N6;1_0 j(=0h.9gjG /Q:%z9:_kI
Fa<Z*Oike[ \93W4̃>x5u ?=&QplxC-c %_Vd#&B̀vj aEx.0pbK53 C%rGl/|54/
@@ 4.3 Ablation Study
Optimization parameters. Figure 5 and 5 show that during feature extraction, optimizing for $n=100$ epochs, each with $m=1,000$ iterations, offers the optimal performance. Additional epochs and optimization iterations, while incurring additional computational cost, do not significantly improve the detection accuracy.
@@
Detection threshold. Figure 7 and 7 show that the system attains higher accuracy, when it adopts a threshold of three votes for considering an input as an anomaly with text only, and four votes with an added detection model using an additional given photo. Setting a high threshold may result in failing to detect an anomaly, while setting a low one may lead to identifying a normal one as anomaly.
@@
Figure 8: Detection accuracy for different numbers of gibberish.
Figure 9: Detection accuracy for different number of real photos.
@@ 5 Defense and Covert Gibberish Generation
In real-world scenarios, the target models being detected may deploy defense mechanisms to recognize anomalous inputs like gibberish, and provide misleading outputs, causing TUNI to misjudge inclusion of PII. To generate more covert gibberish data, we can create strings resembling normal text, with a few characters replaced by syllables from another language. For instance, the detector can craft query texts, by randomly combining English names with syllables from Arabic medical terminology. One way to do this is to start by prompting LLMs like GPT-3.5-turbo to create lists of common initial and final syllables in English words. These syllable lists are then extracted and refined to ensure diversity and eliminate duplicates. Next, the refined syllable combinations are randomly paired to create pseudo-English names, such as “Karinix”, “Zylogene”, “Glycogenyx”, and “Renotyl”. It’s crucial to verify the novelty of these names by checking against a database of real names to avoid collision. Then by prompting the LLM to generate strings using the refined syllable combinations, covert gibberish strings resembling real names are produced (some examples are given in Table 4).
@@
Karinix Zylogene Glycogenyx Renotyl Alphalyn Serinova
Zylotrax Vexilith Dynatrix Solafyne Pyrithex Zephyrax
Exodynix Novylith Glycosyne Thalynex Venatrix Zylocyte
Xenolynx Rynexis Delphylith Myrithex Zetarix Crysolyn
@@ 6 Conclusion
In this paper, we propose TUNI, the first method to conduct identity inference without exposing acutal images to target CLIP models. TUNI turns the inference problem into an anomaly detection problem, through randomly generating textual gibberish that are known to be out of training set, and exploting them to train anomaly detectors. Furthermore, the incorporation of real images is shown to enhance detection performance. Through evaluations across various CLIP model architectures and datasets, we demonstrate the consistent superiority of TUNI over baselines.

### v2

# TUNI: A Textual Unimodal Detector for Identity Inference in CLIP Models

@@ Abstract
The widespread usage of large-scale multimodal models like CLIP has heightened concerns about the leakage of PII. Existing methods for identity inference in CLIP models require querying the model with full PII, including textual descriptions of the person and corresponding images (e.g., the name and the face photo of the person). However, applying images may risk exposing personal information to target models, as the image might not have been previously encountered by the target model. Additionally, previous MIAs train shadow models to mimic the behaviors of the target model, which incurs high computational costs, especially for large CLIP models. To address these challenges, we propose a textual unimodal detector (TUNI) in CLIP models, a novel technique for identity inference that: 1) only utilizes text data to query the target model; and 2) eliminates the need for training shadow models. Extensive experiments of TUNI across various CLIP model architectures and datasets demonstrate its superior performance over baselines, albeit with only text data.
TUNI: A Textual Unimodal Detector for Identity Inference in CLIP Models
Songze Li1,∗,†, Ruoxi Cheng1,thanks: Contributed equally to this work. 1Southeast University, Nanjing China. 2Nanyang Technological University, Singapore. Corresponding authors: songzeli@seu.edu.cn. , Xiaojun Jia2
@@ 1 Introduction
Recent years have witnessed a rapid development of large-scale multimodal models, such as Contrastive Language–Image Pre-training (CLIP) (Radford et al., 2021). These models synthesize information across different modalities, particularly text and images, facilitating applications from automated image generation to sophisticated visual question answering systems. Despite their potential, these models pose significant privacy risks (Inan et al., 2021; Carlini et al., 2021; Leino and Fredrikson, 2020; Rigaki and Garcia, 2023; Helbling et al., 2023; Rahman et al., 2024; Rahman, 2023) as the vast datasets used for training often contain personally identifiable information (PII) (Schwartz and Solove, 2011; Abadi et al., 2016; Bonawitz et al., 2017), raising concerns (Xi et al., 2024) about PII leakage and misuse (Hu et al., 2023; Yin et al., 2021). Therefore, it is extremely important to develop tools to detect potential PII leakage from CLIP models. Specially, as the first step, we would like to address the identity inference problem, i.e., to determine if the PII of a particular person was used in training of a target CLIP model.
Figure 1: Current methods query LLMs with both text and image, while our goal is to conduct identity inference with only textual data.
@@
To address these limitations, we propose a textual unimodel detector (TUNI) for identity inference in CLIP models, which queries the target model with only text information during inference. Specifically, we first propose a feature extractor, which maps a textual description to a feature vector through image optimization guided by the CLIP model; then, we randomly generate a large amount of textual gibberish, which we know do not match any textual descriptions in the training dataset. As shown in Figure 2, we make the key observation that the feature distributions of textual gibberish and member samples in the training set are well distinguishable.
@@ 2.1 Privacy Leakage in CLIP Models
CLIP model exemplifies modern multimodal innovation by integrating an image encoder and a text encoder into its architecture (Radford et al., 2021). These encoders transform inputs into a shared embedding space, enabling effective measurement of semantic similarity (Ramesh et al., 2022). Despite the significant advances and expansive applicability of CLIP models, the vast and diverse datasets utilized for training such models could potentially include sensitive information, raising concerns about privacy leakage (Hu et al., 2022b). Various inference attacks, including model stealing  (Dziedzic et al., 2022; Liu et al., 2022; Wu et al., 2022), knowledge stealing (Liang et al., 2022), data stealing (He and Zhang, 2021), and membership inference attacks (Liu et al., 2021; Ko et al., 2023), have been developed for CLIP, exposing potential vulnerability in privacy leakage. These privacy concerns underscore the necessity for developing robust defense mechanisms to safeguard sensitive information in CLIP models (Golatkar et al., 2022; Jia et al., 2023; Huang et al., 2023).
@@ 2.2 Personally Identifiable Information and Leakage Issues
Personally Identifiable Information (PII) is defined as any data that can either independently or when combined with other information, identify an individual. Training Large Language Models (LLMs) often utilizes publicly accessible datasets, which may inadvertently contain PII. This elevates the risk of data breaches that could compromise individual privacy and entail severe legal and reputational consequences for the deploying entities (Lukas et al., 2023; Abadi et al., 2016; Bonawitz et al., 2017; Rahman et al., 2020; Shamshad et al., 2023). Various attacks have been developed to reveal PII from LLMs. A method is proposed in (Panda et al., 2024) to steal private information from LLMs via crafting specific queries to GPT-4 that can reveal sensitive data by appending a secret suffix to the generated text; Zhang et al. introduced the ETHICIST method for targeted training data extraction, through loss smoothed soft prompting and calibrated confidence estimation, significantly improving extraction performance on public benchmarks (Zhang et al., 2023); Carlini et al. also studied training data extraction from LLMs, emphasizing the predictive capability of attacks given a prefix (Carlini et al., 2021); ProPILE, proposed in (Kim et al., 2024), probes privacy leakage in LLMs, by assessing the leakage risk of PII included in the publicly available Pile dataset; Inan et al. investigated the risks associated with membership inference attacks using a Reddit dataset, further emphasizing the persistent threat of PII leakage in various data environments (Inan et al., 2021).
@@ 3.1 Problem Setup and Threat Model
Detector’s Knowledge and Capability. The detector can query $M$ and observe the output, including extracted image and text embeddings as well as their matching score, but does not know the model architecture of $M$, the parameter values, or the training algorithms. For the target textual description $t$, depending on the application scenarios, the detector may or may not have actual images corresponding to $t$. Nevertheless, in the case where the detector knows corresponding images, due to privacy concerns, it cannot include them in the queries to $M$. The detector cannot modify $M$ or access its internal state.
@@ 3.2 TUNI: Textual Unimodal Detector for ID Inference
Algorithm 1 CLIP-guided Feature Extraction
@@
8:         $v_{x_{j}}\leftarrow M(x_{j})$triangleright Obtain image embedding from $M$
9:         $x_{j+1}\leftarrow\arg\max_{x_{j}}\frac{v_{t}\cdot v_{x_{j}}}{\|v_{t}\|\,\|v_{x_{j}}\|}$triangleright Update image to maximize cosine similarity
@@
11:     $S_{i}\leftarrow\frac{v_{t}\cdot v_{x_{m}}}{\|v_{t}\|\,\|v_{x_{m}}\|}$triangleright Optimized similarity for epoch $i$
12:     $\mathcal{S}\leftarrow\mathcal{S}\cup\{S_{i}\}$, $\mathcal{V}\leftarrow\mathcal{V}\cup\{v_{x_{m}}\}$
@@
16:$D\leftarrow\sqrt{\frac{1}{n}\sum_{v\in\mathcal{V}}\|v-\bar{v}\|^{2}}$
@@
Specifically, for each test text $t$, the detector is equipped with a set of $c$ real images $\{x_{\mathrm{real}}^{1},x_{\mathrm{real}}^{2},\ldots,x_{\mathrm{real}}^{c}\}$. Similar to the feature extraction process in Algorithm 1, over $k$ epochs with independent initializations, $k$ optimized images $\{x_{\mathrm{opt}}^{1},x_{\mathrm{opt}}^{2},\ldots,x_{\mathrm{opt}}^{k}\}$ for $t$ are obtained under the guidance of the CLIP model. Then, we apply a pretrained feature extraction model $F$ (e.g., DeepFace (Taigman et al., 2014) for face images) to the real and optimized images to obtain real embeddings $\{v_{\mathrm{real}}^{1},v_{\mathrm{real}}^{2},\ldots,v_{\mathrm{real}}^{c}\}$ and optimized embeddings $\{v_{\mathrm{opt}}^{1},v_{\mathrm{opt}}^{2},\ldots,v_{\mathrm{opt}}^{k}\}$. Finally, we compute average pair-wise $\ell_{2}$ distance between real and optimized embeddings, denoted by $R$, over $c\cdot k$ pairs, and use $R$ as an additional feature of the text $t$.
4 Evaluations
4.1 Setup
Figure 4: Samples from the dataset for training CLIP models.
Dataset Construction. The datasets for training and ID inference are constructed from three datasets: LAION-5B (Schuhmann et al., 2022), Conceptual Captions 3M (CC3M) (Changpinyo et al., 2021), and FaceScrub (Kemelmacher-Shlizerman et al., 2016). Specifically, 200 celebrities—100 for training and 100 for validation, with their face photos accompanied by labels containing their names are selected from the FaceScrub dataset; then these data samples are augmented by additional photos of the selected celebrities found in LAION-5B, such that each person has multiple photos; finally these augmented data points are mixed with the CC3M dataset to form the training set of the CLIP model. By doing this, we have the ground truth on which people are in the training set and which are not. In our experiments, we construct two datasets, one with a single photo for each person, and another with 75 photos for each person. Samples of this dataset are shown in Figure 4 and a more detailed description is given in appendix.
Baselines. Current ID inference detection methods for CLIP models typically require detector to query target model with corresponding real images. Most MIAs involve training shadow models and related methods like shadow encoders  (Liu et al., 2021), which can be particularly costly for large-scale multimodal models. We empirically compare the performance of TUNI with the following SOTA inference methods, which both avoid using shadow models, but still require submitting both text and image to the target CLIP model for inference.
  • Identity Inference Attack (IDIA) (Hintersdorf et al., 2022) detects with a list of 1000 names to choose from and 30 real photos for a tested person. In IDIA, the attacker (detector) selects candidate names as prompt templates, and predicts names for each image and prompt. Once the correct name is predicted, it’s inferred that the target individual is in training dataset. We compare IDIA using 3 photos for each test sample with TUNI using only text.
4.2 Results
@@ 4.3 Ablation Study
Figure 9: Detection accuracy for different numbers of gibberish.
Figure 10: Detection accuracy for different number of real photos.
Optimization parameters. Figure 6 and 6 show that during feature extraction, optimizing for $n=100$ epochs, each with $m=1,000$ iterations, offers the optimal performance. Additional epochs and optimization iterations, while incurring additional computational cost, do not significantly improve the detection accuracy.
Detection threshold. Figure 8 and 8 show that the system attains higher accuracy, when it adopts a threshold of three votes for considering an input as an anomaly with text only, and four votes with an added detection model using an additional given photo. Setting a high threshold may result in failing to detect an anomaly, while setting a low one may lead to identifying a normal one as anomaly.
@@
Table 3: Samples of randomly generated gibberish.
+̱7IKXb2Y FR!pnI<5xS euiT_;yw/
je|%5(s=G\_ ?Ŵ<E{Dvmz hqf- =j<q5
#lEZ0yrZ5ig ’2_:6[jiOa X*|<tFx|4/
Fa<Z*Oike[ \93W4̃>x5u ?=&QplxC-c
@@ 5 Defense and Covert Gibberish Generation
In real-world scenarios, target models being detected may deploy defense mechanisms to recognize anomalous inputs like gibberish and provide misleading outputs, causing TUNI to misjudge inclusion of PII.
To generate more covert gibberish data, we can create strings resembling normal text, with a few characters replaced by syllables from another language. For instance, the detector can craft query texts, by randomly combining English names with syllables from Arabic medical terminology. One way to do this is to start by prompting LLMs like GPT-3.5-turbo to create lists of common initial and final syllables in English words. These syllable lists are then extracted and refined to ensure diversity and eliminate duplicates. Next, the refined syllable combinations are randomly paired to create pseudo-English names, such as “Karinix”, “Zylogene”, “Glycogenyx”, and “Renotyl”. It’s crucial to verify the novelty of these names by checking against a database of real names to avoid collision. Then by prompting the LLM to generate strings using the refined syllable combinations, covert gibberish strings resembling real names are produced (some examples are given in Table 4).
@@
Karinix Zylogene Glycogenyx
Zylotrax Vexilith Dynatrix
Exodynix Novylith Glycosyne
Xenolynx Rynexis Delphylith
@@ 6 Conclusion
In this paper, we propose TUNI, the first method to conduct identity inference without exposing acutal images to target CLIP models. TUNI turns inference problem into anomaly detection, through randomly generating textual gibberish that are known to be out of training set, and exploting them to train anomaly detectors. Furthermore, the incorporation of real images is shown to enhance detection performance. Through evaluations across various CLIP model architectures and datasets, we demonstrate the consistent superiority of TUNI over baselines.
7 Limitations
Due to constraints resources, we conducted experiments using the name of the individual as textual descriptions. This approach may not fully encapsulate the complexities and nuances of real-world PII leakage including addresses, phone numbers, and other sensitive information.
8 Ethics and Social Impact
The development of TUNI highlights crucial ethical considerations in identity inference using multimodal models like CLIP. By enabling identity inference with only textual data, TUNI reduces the risks associated with exposing PII through images. This approach not only helps protect individual privacy but also minimizes the potential for misuse in harmful applications. As such technologies evolve, it is essential for researchers to adhere to ethical guidelines and promote transparency, ensuring that advancements in AI prioritize user privacy and foster responsible usage in society.
9 Potential Risks
TUNI aims to bolster privacy by aiding in identity inference and safeguarding personal identifiable information within AI systems. While mindful of the risk of misuse, TUNI should adhere to data regulations and be employed only with explicit consent from involved data subjects, promoting privacy and security in AI practices.